You deposit $1,900 in your bank account. If the bank pays 4% simple interest, how much will you accumulate in your account after 10 years?


  Future value $   


b.

What if the bank pays compound interest (annually)? (Do not round intermediate calculations. Round your answer to 2 decimal places.)


  Compound interest $   


c.

How much of your earnings will be interest on interest? (Round your answer to 2 decimal places.)


  Interest on interest $   
Answer #1

The answer has been presented in the supporting sheet. All the parts has been solved with detailed explanation and format. For detailed answer refer to the supporting sheet.

Answer 3 Part 1) 4 Future value = deposit+((deposit* interest rate)* numner of years) 5 = 1900 + ((1900*4%) *10) 6 = 1900+760